You are here

Beam Delivery News

Primary tabs

TItle Date Excerpt Attached images
Beam delivery update: lid up; cyclotron down 27 June 2014

The cyclotron has been vented and the lid raised to allow the electrical fault in the deflector negative plate to be investigated and fixed. The plan at this point is to lower the lid on Wednesday, July 2 and to resume cyclotron operation on or about Monday, July 7.

Beam delivery update: Cyclotron down tomorrow morning for major repairs 26 June 2014

The cyclotron is currently limping along, delivering 9.8 µA to BL2A and beam as needed to BL1B for PIF. This will continue until 06:00 tomorrow (Friday), at which point the cyclotron will be turned off to begin warming up the cryopanels in preparation for venting later in the morning.

Cyclotron down, time to recovery not clear (updated) 25 June 2014

The cyclotron has been down since about 11:30 this morning and will remain off at least overnight due to a problem with the deflector.

Beam delivery update 24 June 2014

We’re running, with 9.8 µA in BL2A and small amounts of beam as needed in BL1B for proton irradiation. BL2C4 is still offline due to a cooling water leak at the solid target facility (STF). Repairing that leak will require the cyclotron to be off for one to 1-1/2 days, likely next week.

Beam delivery update: Maintenance tomorrow 23 June 2014

The cyclotron’s currently delivering ~115 µA to BL1A and 9.8 µA to BL2A. BL2C4 has been off since shortly after 12:00 because of a leak in the STF cooling water system. There is maintenance planned for tomorrow.

Beam delivery update: weekend edition 20 June 2014

We’re currently delivering ~115 µA to BL1A, 9.8 µA to BL2A, and ~95 µA to BL2C4. That is scheduled to continue until Tuesday’s if-needed maintenance.

Beam delivery update 19 June 2014

The cyclotron’s delivering ~115 µA to BL1A, 9.8 µA to BL2A, and ~90 µA to BL2C4. At ISAC, the effort to deliver 95Sr16+ to TIGRESS for experiment S1389 is underway.

Beam delivery update 18 June 2014

The cyclotron’s running relatively well and delivering ~115 µA to BL1A, 8 µA to BL2A, and ~95 µA to BL2C4. At some point the BL2A current may be pushed to 9.8 µA for RIB production at ISAC, but otherwise we expect to run like this for the time being.

Beam delivery update 17 June 2014

We’re currently running with ~110 µA in BL1A, ~2 µA in BL2A, and ~15 µA in BL2C4. The plan for overnight is to ramp up the BL2C4 current and, if RIB tuning at ISAC allows, to carry out a BL2A target scan in advance of increasing the 2A current tomorrow.

Beam delivery update: Long maintenance underway 16 June 2014

This week’s long maintenance period is underway. The Cryogenics/Vacuum group will work to restore normal vacuum in the cyclotron this evening and the cyclotron RF will be restored tomorrow. We still hop to resume operation and beam delivery late afternoon rather than tomorrow evening.

Pages

Subscribe to Group News